OpenStudy (anonymous):

Solve the system by elimination: x + 5y - 4z = -10 2x - y + 5z = -9 2x - 10y - 5z = 0

OpenStudy (jdoe0001):

let us pick 2 off the 3 functions there, and cancel out , say "z" x + 5y - 4z = -10 <---- this one 2x - y + 5z = -9 <---- and this one 2x - 10y - 5z = 0 \(\large \begin{array}{llll} x + 5y - 4z = -10&\times 5\implies & 5x + 25y - 20z = -50\\ 2x - y + 5z = -9&\times 4\implies & 8x - 4y + 20z = -36\\ \hline\\ \textit{sum them up}&\implies&13x-21y+0z=-86\\ \quad \\ \color{blue}{13x-21y=-86} \end{array}\)

OpenStudy (jdoe0001):

let us pick ANOTHER 2 functions off the 3, and again, cancel out "z" x + 5y - 4z = -10 2x - y + 5z = -9 <---- this one 2x - 10y - 5z = 0 <---- and this one \(\large \begin{array}{llll} 2x - y + 5z = -9&\implies & 2x - y + 5z = -9\\ 2x - 10y - 5z = 0&\implies & 2x - 10y - 5z = 0\\ \hline\\ \textit{sum them up}&\implies&4x-11y+0z=-9\\ \quad \\ \color{blue}{4x-11y=-9} \end{array}\) now notice, you have obtained, 2 RESULTANT functions, with only 2 variables, "x" and "y" in the same manner, eliminate either of the variables, and get the other variable

OpenStudy (jdoe0001):

so use the 2 RESULTANT functions to get the value of one variable 13x - 2y = -86 4x -11y = -9

OpenStudy (anonymous):

Thank you this is so helpful!!

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

25y - 4y = ... 21y positive... .ok... so.... minor sum issue there... \(\large \begin{array}{llll} x + 5y - 4z = -10&\times 5\implies & 5x + 25y - 20z = -50\\ 2x - y + 5z = -9&\times 4\implies & 8x + 4y + 20z = -36\\ \hline\\ \textit{sum them up}&\implies&13x+21y+0z=-86\\ \quad \\ \color{blue}{13x+21y=-86} \end{array}\)

OpenStudy (jdoe0001):

well.. \(\large \begin{array}{llll} x + 5y - 4z = -10&\times 5\implies & 5x + 25y - 20z = -50\\ 2x - y + 5z = -9&\times 4\implies & 8x - 4y + 20z = -36\\ \hline\\ \textit{sum them up}&\implies&13x+21y+0z=-86\\ \quad \\ \color{blue}{13x+21y=-86} \end{array}\) rather

OpenStudy (jdoe0001):

so the 2 functions will be 13x + 21y = -86 4x -11y = -9 rechecking... I noticed ... should have been +21y

OpenStudy (anonymous):

Oh okay thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

so z=0, right?

OpenStudy (jdoe0001):

yes, reason why I multiplied by 5 and 4, to have them be -20z and +20z and thus -20z+20z = 0, thus effectively cancelling it out, or ELIMINATING it

OpenStudy (jdoe0001):

notice in the 2nd paiir, I didn't have to do that, because I had a +5z and -5z, and they cancel each other right off

OpenStudy (anonymous):

Okay yeah I was just making sure, thanks!!

OpenStudy (jdoe0001):

np